# An Empirical Study on Challenges for LLM Application Developers

### v3

@@ 7.2.2. External Validity
Limitations of OpenAI LLMs: The threat is that focusing primarily on OpenAI’s LLMs in our empirical study may limit the generalizability of our findings to other LLM vendors. First, by analyzing the number of followers on GitHub accounts for LLMs, we found that ChatGPT has 86.6k followers, while Llama and Gemini have only 5.2k and 1k followers, respectively. Therefore, it can be inferred that the LLMs provided by OpenAI are currently more popular. Second, although different LLM vendors may offer varying features (such as image or audio processing capabilities), we believe that many of the challenges identified in our study (such as API usage, performance optimization, and security issues) are common across different LLMs. This implies that the insights from our study can inform development practices beyond OpenAI’s ecosystem. Finally, to demonstrate the generalizability of our constructed taxonomy, we further analyze challenges related to two other popular LLMs (such as Llama and Gemini) by mining issues from relevant projects on GitHub.
@@ Acknowledgements.
Xiang Chen and Chaoyang Gao have contributed equally to this work and they are co-first authors. Xiang Chen is the corresponding author. This research was partially supported by the National Natural Science Foundation of China (Grant no. 61202006) and the Postgraduate Research & Practice Innovation Program of Jiangsu Province (Grant no. SJCX24_2022).

### v4

@@ 8.2.2. External Validity
Limitations of OpenAI LLMs: The threat is that focusing primarily on OpenAI’s LLMs in our empirical study may limit the generalization of our findings to other LLM vendors. First, by analyzing the number of followers on GitHub accounts for LLMs, we found that ChatGPT has 86.6k followers, while Llama and Gemini have only 5.2k and 1k followers, respectively. Therefore, it can be inferred that the LLMs provided by OpenAI are currently more popular. Second, although different LLM vendors may offer varying features (such as image or audio processing capabilities), we believe that many of the challenges identified in our study (such as API usage, performance optimization, and security issues) are common across different LLMs. This implies that the insights from our study can inform development practices beyond OpenAI’s ecosystem. Finally, to demonstrate the generalizability of our constructed taxonomy, we further analyze challenges related to two other popular LLMs (such as Llama and Gemini) by mining issues from relevant projects on GitHub.
@@
8.2.3. Construct Validity
Determining the popularity of categories: In our empirical study, we use the number of new posts as a proxy for measuring the popularity of categories. However, this metric may not fully reflect true popularity. For instance, older posts tend to accumulate more views over time, potentially introducing bias and making it challenging to accurately evaluate the popularity of newer posts or categories. An alternative approach could involve using the number of views as a metric to measure category popularity. However, the forum only provides cumulative view counts for each post, without the ability to differentiate views by year. As a result, this limitation prevents us from adopting the view-based metric for our analysis.
@@ Acknowledgements.
The authors would like to thank the editor and three anonymous reviewers for their insightful comments and suggestions, which can substantially improve the quality of this work. Xiang Chen and Chaoyang Gao have contributed equally to this work and they are co-first authors. Xiang Chen and Yong Liu are the corresponding authors. This research was partially supported by the National Natural Science Foundation of China (Grant no. 61202006) and the Postgraduate Research & Practice Innovation Program of Jiangsu Province (Grant no. SJCX24_2022).